The transition temperature of a BCS superconductor with an average atomic mass of 199.5 is 4.18K. What is the transition

temperature for an average atomic mass of 203.4?


1
Expert's answer
2021-10-14T11:32:03-0400

The transition temperature is inversely proportional to the square root of the atomic mass:

"T_1=\frac k{\sqrt{m_{a1}}},\\\space\\
T_2=\frac k{\sqrt{m_{a2}}},\\\space\\
\frac{T_1}{T_2}=\sqrt{\frac{m_{a2}}{m_{a1}}},\\\space\\
T_2=T_1\sqrt{\frac{m_{a1}}{m_{a2}}}=4.14\text{ K}."